OpenStudy (anonymous):

collect like terms 6x-9=

OpenStudy (anonymous):

collect like terms 6x-9=

OpenStudy (anonymous):

checking the ans collect like terms 6x-9= 3(2x-3) answer

OpenStudy (anonymous):

that is a true statement. collect like term usually means combining like terms such as 3x+6+7x+3x-2 and turning it into 13x+4

OpenStudy (anonymous):

so mrjoselson my answer is correct

OpenStudy (anonymous):

yes, you factored the original express to get an equivalent expression